# Moduli spaces and breather dynamics of analytic solutions in chiral magnets without Heisenberg exchange interaction

### v1

@@ 2.3 Infinite-dimensional family of non-axisymmetric solutions
By the exact same logic, we can find an infinite-dimensional family of solutions, by allowing the centre of the circle at each radius to be different. Reintroducing the centre $\vec{X}_{c}$, any function $\vec{X}_{c}(R)$ between $R=0$ and $R=R_{\text{max}}$ will give a valid solution (granted that $|\vec{X}_{c}^{\prime}(R)|<1$ as we will see shortly). Semi-explicitly, we can write this solution in polar-like co-ordinates adapted to this function:
@@
where $\varphi(\vec{X}_{c})$ means azimuthal co-ordinates with $\vec{X}_{c}$ as the centre, and $R^{2}=(x_{1}-X_{c,1}(R))^{2}+(x_{2}-X_{c,2}(R))^{2}$. This equation has $R$ on both sides and is thus hard to effect as a co-ordinate transformation for a general function.
In order for circles of constant $n_{3}$ not to intersect (which will relate to the well-definedness of this co-ordinate transformation), this function must satisfy
$\lvert\vec{X}_{c}^{\prime}(R)\rvert<1\ ,\quad\forall R\in[0,R_{\text{max}}]\ .$ (24)
This gives a moduli space of static Skyrmion-like solutions described by the space
$\{\vec{X}_{c}\in C_{0}([0,R_{\text{max}}],\mathbb{R}^{2}):\lvert\vec{X}_{c}^{\prime}(R)\rvert<1\text{ where it is defined}\}\ .$ (25)
The simplest solution, $\vec{X}_{c}^{\prime}(R)=\vec{0}$, gives the axisymmetric solutions above, while in general a solution will still be contained within a circle of radius $R_{\text{max}}$ but have circles of constant $n_{3}$ at different displacements within. “Where it is defined” allows for the curve $\vec{X}_{c}(R)$ to have corners.
The boundaries of this moduli space are given by $\lvert\vec{X}_{c}^{\prime}(R)\rvert=1$ for some $R$. In the extreme case where $\vec{X}_{c}^{\prime}(R)=\vec{e}_{1}$ for all $R$, the centre of the Skyrmion touches the edge and at this point the magnetic field is not defined. In appendix A, a sanity check is presented to verify the validity of the fluid flow equations evaluated on the solutions above.
2.4 Infinite-dimensional symmetry of the energy
We check the result in a different way, that also reveals the more general symmetry of the model. We parametrize the map $\bm{n}=(\vec{\nu},n_{3})$ with $\vec{\nu}=(n_{1},n_{2})$ and $n_{3}=\sqrt{1-\nu_{i}\nu_{i}}$. Thus the energy for the restricted model can be written as
$\begin{array}[]{rlc}E(\bm{n})&=\displaystyle\int\mathop{}\!\mathrm{d}^{2}x\Big{(}2kn_{3}\epsilon_{ij}\partial_{i}\nu_{j}+V(n_{3})\Big{)}\\ &\equiv\displaystyle E_{1}+E_{0}=\int\mathop{}\!\mathrm{d}^{2}x\big{(}{\cal E}_{1}+{\cal E}_{0}\big{)}\ ,\end{array}$ (26)
where we have integrated by parts. Let us take a generic map $x_{i}\to\bm{n}$, not necessarily a solution of the equation of motion. Then consider $\gamma(n_{3})$, the preimage of $n_{3}$ constant. Let us call $a(\gamma(n_{3}))$ the area inside the curve $\gamma(n_{3})$. We then can write the potential contribution to the energy as
$E_{0}=\int\mathop{}\!\mathrm{d}n_{3}\frac{\mathop{}\!\mathrm{d}a}{\mathop{}\!\mathrm{d}n_{3}}V(n_{3})\ .$ (27)
We notice that all the maps with the same area $a(\gamma(n_{3}))$ also have the same $E_{0}$. In particular this is achieved by all maps $\Gamma:\mathbb{R}^{2}\to\mathbb{R}^{2}$ where $\gamma(n_{3})$ and $(F\circ\gamma)(n_{3})$ are connected by a rigid translation. Notice that $\Gamma$ is not necessarily an area-preserving diffeomorphism. So ${\cal E}_{0}$ is not necessarily invariant, but $E_{0}$ is invariant. Let us define the circulation of $\vec{\nu}$ around the curve $\gamma$,
$c(n_{3})=\oint_{\gamma(n_{3})}\mathop{}\!\mathrm{d}\vec{\ell}\cdot\vec{\nu}\ ,$ (28)
and by using Stokes theorem we can write the DMI term as
$E_{1}=\int kn_{3}\frac{\mathop{}\!\mathrm{d}c(n_{3})}{\mathop{}\!\mathrm{d}n_{3}}\mathop{}\!\mathrm{d}n_{3}\ .$ (29)
$E_{1}$ remains invariant if and only if $c(n_{3})$ is invariant. This is the case of rigid translations of $\gamma(n_{3})$. So the maps where $\gamma(n_{3})$ and $(\Gamma\circ\gamma)(n_{3})$ are connected by a rigid translation, are symmetries of both functionals $E_{0}$ and $E_{1}$, and consequently of $E(\bm{n})$. We may call these symmetries “contour-preserving” maps. Thus we have an $\infty$-dimensional symmetry group, $\mathcal{T}^{\rm rigid}_{2}$, of the whole energy functional (26), valid for all maps, not only the solutions of the equations of motion.
We can thus calculate the stress-energy tensor through Noether’s theorem. The total Lagrangian density $\mathcal{L}$ for this system, assuming a static configuration, is just the negative of the energy density, $\mathcal{L}=-(\mathcal{E}_{1}+V(n_{3}))$. The energy-momentum tensor associated to rigid translationally invariant symmetry, is given by
$\begin{array}[]{rlc}T_{ab}&=\displaystyle-\frac{\partial\mathcal{L}}{\partial(\partial^{a}\nu_{i})}\partial_{b}\nu_{i}+\delta_{ab}\mathcal{L}\\ \\ &=\displaystyle 2kn_{3}\epsilon_{aj}\partial_{b}\nu_{j}-\delta_{ab}(2kn_{3}\epsilon_{ck}\partial_{c}\nu_{k}+V(n_{3}))\ ,\end{array}$ (30)
where $a,b,c=1,2$ are the spatial indices. This model is not fully invariant under spatial translations due to the presence of the external potential term $V(n_{3})$. However, the model is invariant under rigid (global) translations, which means translations that shift the entire configuration of $\bm{n}$ uniformly without altering its structure relative to the co-ordinates, i.e. $\bm{n}(\vec{x})\to\bm{n}(\vec{x}-\vec{a})$.
Being $n_{3}$ fixed and constant, the system reduces effectively to a two-dimensional vector field $\nu$, subject to the constraint $\nu_{i}\nu_{i}=1-n_{3}^{2}=\mathrm{const.}$, leading to an $S^{1}$-symmetry in the $(n_{1},n_{2})$-plane. Finally, the total infinite-dimensional symmetry group of the restricted model is given by
$\mathcal{G}_{\infty}=\mathcal{T}^{\rm rigid}_{2}\ltimes\operatorname{O}(2)\ .$ (31)
@@ 2.5 Effective string-like model
As we noticed above in Eq. (25), the moduli space can be parametrized by $\vec{X}_{c}(R)$, the line of the centres depending on $R(n_{3})$ as in Eq. (13). From this interesting prospective, the moduli space is a sort of “string model” with $R$ parametrizing the string and $\vec{X}_{c}$ giving its embedding in space. This is shown pictorially in Fig. 3. In the chiral magnet model without the Heisenberg term, all the configurations (24) have the same energy; thus the effective string would have an infinite square-well elastic potential, tensionless up to some maximum extension. At first level of approximation, we can compute the effect of the Heisenberg contribution by evaluating $E_{2}$ on the solution of the model (26). We know that only the $\operatorname{SO}(2)$ invariant solutions with $\vec{X}^{\prime}_{c}(R)=\vec{0}$ will be the true minima of the energy.
@@ 2.6 Explicit solutions
We will now construct a few examples of explicit solutions and numerically verify that they indeed satisfy the full equations of motion. The solutions are shown in Fig. 4.
@@
where we have chosen $h=1$, $u=0$ leading to supercompactons with $R\in[0,2]$. The evaluation of the equation of motion (4) vanishes to numerical precision of the derivatives computed numerically on an interpolated field in the Mathematica package. This confirms that the collection of fluid paths indeed gives a full solution to the linear PDE.
@@ 3 Dynamics of breather-like solutions
We can also understand dynamical solutions through these characteristics, by relaxing the condition that flows are stationary. Note that from (10) we can define $\omega(n_{3})=-f^{\prime}(n_{3})$, which represents the frequency with which a single spin, followed along its trajectory, returns to the initial point.
@@
Systematically, we should consider a set of characteristics with equal $n_{3}$, parametrized by periodic variable $\bar{\sigma}=\frac{\sigma}{2\pi R(n_{3})}$, starting tangent to a circle of radius $R(n_{3})+\Delta(n_{3})$ at time $\Phi(n_{3})/\omega(n_{3})$ (that is, we allow different oscillations to have different phases $\Phi$). For simplicity we take $R>0$ and thus $R^{\prime}(n_{3})>0$ in this section; other possibilities are related by symmetry. Note that $\Delta$, which we can also take as positive without loss of generality, is the initial perturbation profile, it is not necessarily small. As before, we have complex co-ordinates $X$ giving the position of the characteristic and $\nu$ giving the tangent direction and speed of the characteristic with respect to time. Since we are now considering a one-parameter family of characteristics, $X$ and $\nu$ are functions of two variables, $t$ and $\bar{\sigma}$:
@@
$\mathcal{R}=\sqrt{R(n_{3})^{2}+\Delta(n_{3})^{2}+2R(n_{3})\Delta(n_{3})\cos(\omega(n_{3})t-\Phi(n_{3}))}\ .$ (39)
Note that to avoid this $\mathcal{R}$-circle to not collapse to zero size, we require $\lvert\Delta\rvert<R$. To construct an extended solution, we want to put together the circles of different constant $n_{3}$ inside each other. In general, this is not possible since the period of oscillation $\frac{2\pi}{\omega(n_{3})}$ will be different for each $n_{3}$, and thus we cannot avoid the different circles colliding, at which point our field becomes discontinuous. However, with pure Zeeman interaction, that is the potential (7) with $u=0$, $\omega(n_{3})=h$ and thus all circles have the same period of oscillation. Therefore, to look for breathing solutions, we must specialize to a model with only Zeeman interaction in the potential.
Even if all circles oscillate in radius at the same time, they have different amplitudes that depend on their initial radius, so we must choose those initial radii so that the circles never collide, or shrink to zero radius. Mathematically, this means finding functions $\Delta(n_{3})$ and $\Phi(n_{3})$ such that
$\begin{cases}\Delta(-1)=0\\ \Delta<R\\ \partial_{n_{3}}\mathcal{R}^{2}>0\ ,\quad\quad 0<t<\frac{2\pi}{h}\ ,\end{cases}$ (40)
for $n_{3}$ up to some maximum which may or may not be the same as $n_{3}(R_{\text{max}})$.
@@
$\Delta\Delta^{\prime}+RR^{\prime}+(R\Delta^{\prime}+R^{\prime}\Delta)\cos(ht-\Phi)-R\Delta\Phi^{\prime}\sin(ht-\Phi)>0\ .$ (41)
@@
$\Delta\Delta^{\prime}+RR^{\prime}-\sqrt{(R\Delta^{\prime}+R^{\prime}\Delta)^{2}+(R\Delta\Phi^{\prime})^{2}}>0\ ,$ (42)
which we can rearrange to
$(R^{\prime 2}-\Delta^{\prime 2})(R^{2}-\Delta^{2})-R^{2}\Delta^{2}\Phi^{\prime 2}>0\ .$ (43)
We see that since $\Delta^{2}<R^{2}$, it is a necessary condition that $\Delta^{\prime 2}<R^{\prime 2}$. This then provides a constraint on the maximum size of $\Phi^{\prime}$ wherever $\Delta\neq 0$ (where it is equal to zero $\Phi$ is undefined).
This gives an infinite-dimensional moduli space of periodic axisymmetric solutions. These resemble the breathing Skyrmion solutions observed in numerics, where the radius oscillates and is coupled to rotations of the helicity, that is the angle $\delta$ in Eq. (44).
@@
Figure 5: Constant $n_{3}$ contours at different points in time. Note that while the magnetization is always tangent to its path-line, for $ht-\Phi\neq n\pi$ it is not tangent to the circle of constant $n_{3}$, in contrast to the axisymmetric static solution.
Note that it is valid, within these constraints, to have a $\Delta$ that is supported only in a subset of $[-1,n_{3}(R_{\text{max}})]$. Physically, this corresponds to a Skyrmion that only “breathes” within a certain annulus, and it is otherwise static.
In contrast to the static case, where attainment of the inequality meant an invalid solution, here $\lvert\Delta\rvert=R$ or $\lvert\Delta^{\prime}\rvert=R^{\prime}$ at some $n_{3}$ will in general give the development of a singularity in finite time. This could occur at a point, at the centre of the Skyrmion, or along a circle.
@@
Note that the $\operatorname{U}(1)$ action where we rotate both $\bm{n}$ and $\vec{x}$ in opposite directions is a symmetry of the energy, and this symmetry is preserved by Landau-Lifshitz dynamics. Thus we would be justified in constraining the dynamical Euler-Lagrange equations to the most general $\operatorname{U}(1)$ ansatz:
$\bm{n}=\begin{pmatrix}\sin\Theta(r)\cos(\phi+\delta(r))\\ \sin\Theta(r)\sin(\phi+\delta(r))\\ \cos\Theta(r)\end{pmatrix}\ ,$ (44)
where $(r,\phi)$ are polar co-ordinates in the plane. If $\delta(r)$ is constant, it is the helicity variable described above. Solving the resulting coupled ODEs for $\Theta(r)$ and $\delta(r)$ should then give the same set of solutions.
@@
This reflects the coupling between helicity and breathing that has been observed in collective co-ordinate approximations for breathing magnetic Skyrmions [22], where an ansatz is assumed with $\delta$ independent of $r$. We can find special solutions with such $\delta$ when $\Delta/R$ and $\Phi$ are constants independent of $n_{3}$.
@@
$RR^{\prime}+\Delta\Delta^{\prime}+(R^{\prime}\Delta+R\Delta^{\prime})\cos(ht-\Phi)+R\Delta\Phi^{\prime}\sin(ht-\Phi)$ (49)
-$\mathop{-}\sqrt{R^{2}+\Delta^{2}+2R\Delta\cos(ht-\Phi)}\lvert X_{c}^{\prime}\rvert>0\ .$

### v2

@@ 2.3 Infinite-dimensional family of non-axisymmetric solutions
In the section above, we impose an unnecessary constraint by asking for axisymmetry, which equivalently means that the circular path lines are concentric. We can imagine that by varying the centres of the path lines as a function of the radius, we can construct a large family of solutions with the same energy and topological charge as the axisymmetric one. In this section we ask what is the most general continuous solution we can construct from closed path lines, in general on subsets of the plane bounded by the largest possible path line, if it exists. Such solutions can be put together to make a subset of all discontinuous solutions in the plane.
Trivially, we can construct the uniform state $n_{3}=1$ or $n_{3}=-1$ on any domain, since each path line is closed. If we want to construct a solution which differs from the uniform state, it must have path lines that are circles. If the path lines can have arbitrarily large radius, we can again construct a solution in the whole plane. If instead there is a largest possible radius $R_{\text{max}}$, the largest region on which we can define such a solution is the disc $D$ of radius $R_{\text{max}}$, $B(R_{\text{max}})$, with boundary conditions that $\vec{\nu}(\partial B(R_{\text{max}}))$ is parallel to the boundary. The axisymmetric solution above is one non-uniform example that satisfies the boundary conditions. Here we address the question of what other non-uniform solutions exist on this disc.
For now we assume that $h\neq\pm 2u$ so that $n_{3}^{\star}\neq\pm 1$, so that $n_{3}=n_{3}^{\star}$ is a circle dividing the target sphere into two regions. Any continuous solution must attain all values on the target space in one of the two regions on the target sphere bounded by $n_{3}=n_{3}^{\star}$. To disambiguate which region, we specify if $n_{3}$ attains +1+1 or 1-1, meaning it must attain all values in the set $(n_{3}^{\star},1]$ or $1,n_{3}^{\star})$ respectively. This motivates the definition:
Definition 1.
The moduli spaces $\mathcal{M}^{+}_{h,u}$, $\mathcal{M}^{-}_{h,u}$, are the spaces of all continuous and differentiable solutions to the Euler-Lagrange equations with the standard potential (7),
  1. (i)
    on the disc with radius $R_{\rm max}(h,u)$ (18) satisfying the boundary conditions $n_{3}=n_{3}^{*}(h,u)$ and $\vec{\nu}$ tangent to the boundary, if $R_{\rm max}(h,u)$ is finite
  2. (ii)
    on the plane, satisfying the boundary conditions at infinity $n_{3}\to n_{3}^{*}(h,u)$, $\vec{\nu}\cdot\hat{\vec{r}}\to 0$, if $R_{\rm max}(h,u)$ is infinite
attaining $n_{3}=\pm 1$ respectively in the interior.
These definitions of the moduli spaces also make sense for the special cases $h=\pm 2u$, but $\mathcal{M}^{+}_{-2u,u}$ is empty, as is $\mathcal{M}^{-}_{2u,u}$. For $0\leq h\leq 2u$, the moduli spaces formally exist but the solutions they describe are unstable, in the same way as described in Sec. 2.2.
Since such solutions are relative homotopies of the axisymmetric solutions, the topological charge of all solutions in $\mathcal{M}_{u,h}^{\pm}$ is the same as in (21).
Note that the boundary condition in case (ii) breaks translation symmetry by picking a preferred origin that lies ‘at the centre of the path line at infinity’, so a fuller moduli space of continuous plane-filling solutions can be found by taking $\mathcal{M}_{h,u}^{\pm}\times\mathbb{R}^{2}$.
With this definition, we can equate valid static solutions of the Euler-Lagrange equations to a certain function space:
Theorem 2.
For $h\neq\pm 2u$, $\mathcal{M}^{+}_{h,u}$ and $\mathcal{M}^{-}_{h,u}$ are in bijection with each other and with the space of $C^{1}$ functions, $\vec{X}_{c}$, from:
  1. (i)
    the closed interval $0,R_{\text{max}}(h,u)]$ to the disc $B(R_{\text{max}})$, with boundary condition $\vec{X}_{c}(R_{\text{max}})=\vec{0}$, if $R_{\text{max}}(h,u)$ is finite,
  2. (ii)
    the half-line $0,\infty)$ to the plane, with boundary condition $\lim_{R\to\infty}\vec{X}_{c}(R)=\vec{0}$, if $R_{\text{max}}(h,u)$ is infinite,
whose gradient has a norm always smaller than 1.
Informally, the proof follows from the fact that contours of constant $n_{3}$ of any static continuous solution form a kind of foliation of the domain by circles and a point, which can be parametrized by the positions of the centres of the circles of each radius. We call the function that takes in a radius and outputs the centre of the $n_{3}$ contour of that radius $\vec{X}_{c}(R)$, and find that it must belong to the above function space, and conversely that any function belonging to that function space generates a valid solution.
Formally, first we show that any $\bm{n}$ belonging to $\mathcal{M}_{h,u}^{\pm}$ is described by a function $\vec{X}_{c}$ that is in the above function space.
Firstly, the boundary conditions and continuity ensure that $\bm{n}$ attains all values in one of the two spherical caps bounded by $n_{3}=n_{3}^{\star}$. We can consider the preimage of a set of constant $n_{3}$ in the interval $(-1,n_{3}^{\star})$ or $(n_{3}^{\star},1)$ for $\bm{n}$ in the moduli space $\mathcal{M}^{-}_{h,u}$, $\mathcal{M}^{+}_{h,u}$ respectively. It must then be a circle or set of circles inside the domain, of given radius $R(n_{3})$ as in (12). So to a given $n_{3}$ and thus a given $R\in[0,R_{\text{max}})$ (since $R(n_{3})$ is invertible), we have in general a set of points $\{\vec{X}_{c}^{1},\vec{X}_{c}^{2}\ldots\}$ that are the centres of the these circles, as in the general path line solution (10).
To form a valid solution, these circles (and points, for $R=0$) must foliate the plane, i.e. not intersect and also fill every point. By definition, these circles fill the plane. Non-intersection of circles means that one of two inequalities is satisfied for two distinct circles:
$\begin{cases}\lvert\vec{X}_{c}^{i}(R^{A})-\vec{X}_{c}^{j}(R^{B})\rvert<\lvert R^{A}-R^{B}\rvert&\text{circles nested}\\ \lvert\vec{X}^{i}_{c}(R^{A})-\vec{X}^{j}_{c}(R^{B})\rvert>R^{A}+R^{B}&\text{otherwise}\end{cases}$ (22)
The key claim we must prove is that any point outside a circular path line lies on a circular path line encircling the first path line. Thus, all circles are nested, and to each $R$ we can associate a single $\vec{X}_{c}(R)$.
To see this, first consider an open neighbourhood $U(\vec{x})$ of a point $\vec{x}$ on a circular path line $C(\vec{x})$ which is not a point, i.e. $R(n_{3}(\vec{x}))\neq 0$. The curve $C(\vec{x})$ divides $U(\vec{x})$ in two. Let us call the outward normal to the path line $\vec{e}(C(\vec{x}))$. For any point $\vec{x}^{\prime}$ in $U$ such that $(\vec{x}^{\prime}-\vec{x})\cdot\vec{e}\geq 0$, the corresponding path line $C(\vec{x}^{\prime})$ must enclose $C(\vec{x})$ (and thus $R(n_{3}(\vec{x}^{\prime}))>R(n_{3}(\vec{x}))$), as the latter inequality in (22) cannot be satisfied for sufficiently small neighbourhood.
Let us now consider any $\vec{x}_{1}$ in the plane outside $C(\vec{x})$. We can define $\vec{x}_{0}$ as the closest point of $C(\vec{x})$ to $\vec{x}_{1}$, and define a straight path interpolating between them parametrized by $\sigma$, $\vec{x}(\sigma)=\vec{x}_{0}+\sigma(\vec{x}_{1}-\vec{x}_{0})$.
Now we consider the scalar function $\vec{e}(C(\vec{x}(\sigma)))\cdot(\vec{x}_{1}-\vec{x}_{0})$. It must remain positive: to see this, suppose that there existed some $\vec{x}_{2}$ such that $\vec{e}(C(\vec{x}_{2}))\cdot(\vec{x}_{1}-\vec{x}_{0})=0$. Considering a point $\vec{x}_{3}$ between $\vec{x}_{0}$ and $\vec{x}_{2}$ in the neighbourhood of $\vec{x}_{2}$, then $(\vec{x}_{3}-\vec{x}_{2})\cdot\vec{e}=0$ therefore $C(\vec{x}_{3})$ encloses $C(\vec{x}_{2})$ and thus $\vec{e}(C(\vec{x}_{3}))\cdot(\vec{x}_{1}-\vec{x}_{0})<0$. That is, the scalar function defined above must be negative in the neighbourhood immediately before a zero, and thus cannot decrease from positive to zero as $\sigma$ is increased.
Therefore, $R$ increases continuously across this line and $C(\vec{x}^{\prime})$ encloses $C(\vec{x})$. Having established that all contours form one ordered set of nested circles, we know that there is only a single circle of each radius. We can thus define a single-valued map $\vec{X}_{c}(R)$ that for a given radius $R$ gives the point in the plane that is the centre of the path line with radius $R$, with constant value $n_{3}$ such that $R=R(n_{3})$. For all distinct values of $R$, this function must satisfy the first inequality in (22):
$\lvert\vec{X}_{c}(R^{A})-\vec{X}_{c}(R^{B})\rvert<\lvert R^{A}-R^{B}\rvert\$. (23)
To proceed it is useful to write the solution $\bm{n}$ corresponding to a given $\vec{X}_{c}$ in polar-like co-ordinates adapted to $\vec{X}_{c}$:
@@
where $\varphi(\vec{X}_{c})$ means azimuthal co-ordinates with $\vec{X}_{c}$ as the centre, and $R^{2}=(x_{1}-X_{c,1}(R))^{2}+(x_{2}-X_{c,2}(R))^{2}$. This equation has $R$ on both sides and is thus hard to effect as a co-ordinate transformation for a general function. The constant function $\vec{X}_{c}(R)=\vec{0}$ gives the axisymmetric solutions in Sec. 2.2, while in general a solution will have circles of constant $n_{3}$ at different displacements within.
Since $\vec{X}_{c}(R)$ is Lipschitz, it is differentiable except at isolated points. If we assume that $\vec{X}_{c}(R)$ does not have a derivative at a point, we find that $\bm{n}$ is not differentiable. To see this, we use the fact that locally, such a point of non-differentiability can be represented by
$\vec{X}_{c}(R)=\begin{cases}v_{A}R\vec{e}_{A}&0\leq R\leq R_{\star}\\ v_{A}R_{\star}+v_{B}(R-R_{\star})\vec{e}_{B}&R_{\star}\leq R\leq R_{\text{max}}\end{cases},$ (25)
with $0<R^{\star}<R_{\text{max}}$, $\vec{e}_{A,B}$ arbitrary unit vectors in the plane and $0\leq v_{A,B}<1$. This form is simple enough to find the co-ordinate transform explicitly, and thus compute $\nabla\bm{n}$ and find it is discontinuous unless $v_{A}=v_{B}$ and $\vec{e}_{A}=\vec{e}_{B}$, i.e. if there is no discontinuity in $\vec{X}_{c}^{\prime}(R)$. Since $\vec{X}_{c}(R)$ must be differentiable, equation (23) becomes
$\lvert\vec{X}_{c}^{\prime}(R)\rvert<1\ .$ (26)
In case (i), we have the boundary condition on the map $R_{\text{max}}$ to $\vec{0}$, the origin of the disc. In case (ii), the boundary condition tells us that $\lim_{R\to\infty}\vec{X}_{c}=\vec{0}$. The constraints on $\bm{n}$ thus give us the constraints on $\vec{X}_{c}$ described in Thm. 2.
[…]
$\partial_{\vec{x}}\tilde{r}=\frac{(\vec{x}-\vec{X}_{c})\cdot\partial_{\vec{x}}\vec{x}}{\tilde{r}+(\vec{x}-\vec{X}_{c})\cdot\vec{X}^{\prime}(R)}\ .$ (27)
The denominator of this fraction cannot be zero, since on the one hand $\tilde{r}=\lvert\vec{x}-\vec{X}_{c}\rvert$, and on the other $\lvert(\vec{x}-\vec{X}_{c})\cdot\vec{X}_{c}^{\prime}(R)\rvert\leq\lvert\vec{x}-\vec{X}_{c}\rvert\lvert\vec{X}_{c}^{\prime}(R)\rvert<\lvert\vec{x}-\vec{X}_{c}\rvert$. Therefore $\tilde{r}(\vec{x})$ is differentiable. We thus establish a one-to-one mapping between solutions to the Euler-Lagrange equations and functions in the function space described in Thm. 2. This concludes the proof.
@@ 2.4 Effective string-like model
As we noticed above in Eq. (23), the moduli space can be parametrized by $\vec{X}_{c}(R)$, the line of the centres depending on $R(n_{3})$ as in Eq. (13). From this perspective, the moduli space is a sort of “string model” with $R$ parametrizing the string and $\vec{X}_{c}$ giving its embedding in space. This is shown pictorially in Fig. 3. In the chiral magnet model without the Heisenberg term, all the configurations (23) have the same energy; thus the effective string would have an infinite square-well elastic potential, tension-less up to some maximum extension. At first level of approximation, we can compute the effect of the Heisenberg contribution by evaluating $E_{2}$ on the solutions (24) of the model without Heisenberg interaction. We know that only the $\operatorname{SO}(2)$ invariant solutions with $\vec{X}^{\prime}_{c}(R)=\vec{0}$ will be the true minima of the energy (1).
@@ 2.5 Explicit solutions
We will now illustrate a few examples of explicit solutions. The solutions are shown in Fig. 4.
@@
where we have chosen $h=1$, $u=0$ leading to supercompactons with $R\in[0,2]$. We checked that the evaluation of the equation of motion (4) vanishes to numerical precision of the derivatives computed numerically on an interpolated field in the Mathematica package.
@@ 3 Dynamics of breather-like solutions
We can also understand dynamical solutions through these characteristics, by relaxing the condition that flows are stationary. Note that from (10) we can define $\omega(n_{3})=-U^{\prime}(n_{3})$, which represents the frequency with which a single spin, followed along its trajectory, returns to the initial point.
@@
Systematically, we should consider a set of characteristics with equal $n_{3}$, parametrized by periodic variable $\bar{\sigma}=\frac{\sigma}{2\pi R(n_{3})}$, starting tangent to a circle of radius $R(n_{3})+\Delta(n_{3})$ at time $\Phi(n_{3})/\omega(n_{3})$ (that is, we allow different oscillations to have different phases $\Phi$). We can take $\Delta>0$ without loss of generality. Note that $\Delta$ is the initial perturbation profile, it is not necessarily small. As before, we have complex co-ordinates $X$ giving the position of the characteristic and $\nu$ giving the tangent direction and speed of the characteristic with respect to time. Since we are now considering a one-parameter family of characteristics, $X$ and $\nu$ are functions of two variables, $t$ and $\bar{\sigma}$:
@@
$\mathcal{R}(n_{3},t)=\sqrt{R(n_{3})^{2}+\Delta(n_{3})^{2}+2R(n_{3})\Delta(n_{3})\cos(\omega(n_{3})t-\Phi(n_{3}))}\ .$ (35)
To avoid this $\mathcal{R}$-circle to not collapse to zero size, we require $\Delta<\lvert R\rvert$.
To construct an extended solution, we want to put together the circles of different constant $n_{3}$ inside each other. In general, this is not possible since the period of oscillation $\frac{2\pi}{\omega(n_{3})}$ will be different for each $n_{3}$, and thus we cannot avoid the different circles colliding, at which point our field becomes discontinuous. However, with pure Zeeman interaction, that is the potential (7) with $u=0$, $\omega(n_{3})=h$ and thus all circles have the same period of oscillation. Therefore, to look for breathing solutions, we must specialize to a model with only Zeeman interaction in the potential.
Even if all circles oscillate in radius at the same time, they have different amplitudes that depend on their initial radius, so we must choose those initial radii so that the circles never collide, or shrink to zero radius.
This means finding $\Delta(n_{3})$, $\Phi(n_{3})$ such that
$\begin{cases}\Delta(-1)=0\\ \Delta<R\\ \partial_{n_{3}}\mathcal{R}^{2}>0\ ,\quad\quad 0<t<\frac{2\pi}{h}\ .\end{cases}$ (36)
@@
$\Delta\Delta^{\prime}(n_{3})+RR^{\prime}(n_{3})+(R\Delta^{\prime}(n_{3})+\Delta R^{\prime}(n_{3}))\cos(ht-\Phi)-R\Delta\Phi^{\prime}(n_{3})\sin(ht-\Phi)>0\ .$ (37)
@@
$\Delta\Delta^{\prime}(n_{3})+RR^{\prime}(n_{3})-\sqrt{(R\Delta^{\prime}(n_{3})+R^{\prime}(n_{3})\Delta)^{2}+(R\Delta\Phi^{\prime}(n_{3}))^{2}}>0\ .$ (38)
This constrains $\Delta(R)$, $\Phi(R)$ in a similar way to the way that $\vec{X}_{c}(R)$ is constrained in (23). This constraint can be written as
$\lvert\Delta^{\prime}(R)\rvert<1,\,\lvert\Phi^{\prime}(R)\rvert<\sqrt{(1-\Delta^{\prime}(R)^{2})(\lvert\Delta\rvert^{-2}-\lvert R\rvert^{-2})}\ .$ (39)
For an illustration, see Fig. 5.
@@
Figure 5: Constant $n_{3}$ contours at different points in time. Note that while the magnetization is always tangent to its path line, for $ht-\Phi\neq n\pi$ it is not tangent to the circle of constant $n_{3}$, in contrast to the axisymmetric static solution.
Note that it is valid, within these constraints, to have a $\Delta$ that is supported only in a subset of $1,n_{3}^{\star}]$ or $n_{3}^{\star},1]$. Physically, this corresponds to a Skyrmion that only “breathes” within a certain annulus, and it is otherwise static.
In contrast to the static case, where attainment of the inequality meant an invalid solution, here $\Delta=\lvert R\rvert$ or $\lvert\Delta^{\prime}\rvert=\lvert R^{\prime}\rvert$ at some $n_{3}$ will in general give the development of a singularity in finite time. This could occur at a point, at the centre of the Skyrmion, or along a circle.
@@
Note that the $\operatorname{U}(1)$ action where we rotate both $\bm{n}$ and $\vec{x}$ in opposite directions is a symmetry of the energy, and this symmetry is preserved by Landau-Lifshitz dynamics. Thus we can consistently constrain the Landau-Lifshitz equation to the most general $\operatorname{U}(1)$ ansatz:
$\bm{n}=\begin{pmatrix}\sin\Theta(r)\cos(\phi+\delta(r))\\ \sin\Theta(r)\sin(\phi+\delta(r))\\ \cos\Theta(r)\end{pmatrix}\$, (40)
where $(r,\phi)$ are polar co-ordinates in the plane. Solving the resulting coupled ODEs for $\Theta(r)$ and $\delta(r)$ should then give the same set of solutions.
@@
If $\delta(r)$ is constant, it is the ‘Skyrmion helicity’ variable commonly referred to in the literature. This equation reflects the coupling between helicity and breathing that has been observed in collective co-ordinate approximations for breathing magnetic Skyrmions [23], where an ansatz is assumed with $\delta$ independent of $r$. We can find special solutions with such $\delta$ when $\Delta/R$ and $\Phi$ are constants independent of $n_{3}$.
@@
$RR^{\prime}(n_{3})+\Delta\Delta^{\prime}(n_{3})+(R^{\prime}(n_{3})\Delta+R\Delta^{\prime}(n_{3}))\cos(ht-\Phi)+R\Delta\Phi^{\prime}(n_{3})\sin(ht-\Phi)$ (45)
-$\mathop{-}\sqrt{R^{2}+\Delta^{2}+2R\Delta\cos(ht-\Phi)}\lvert X_{c}^{\prime}(n_{3})\rvert>0\ .$